Derive schrodingers equation

1
Expert's answer
2021-12-29T13:06:56-0500

We know that wave equation

"\psi(x,t)=Ae^{i(kx-wt)}"

Hamilton equation

H=T+V

Now

"E=\frac{p^2}{2m}+V(x)"

Take first derivative of a function


"\frac{d\psi(x,t)}{dt}=-iwAe^{(kx-wt)}=-iw\psi(x,t)"

"\frac{d^2\psi(x,t)}{dt^2}=-k^2Ae^{(kx-wt)}=-k^2\psi(x,t)"

"\frac{d^2\psi(x,t)}{dt^2}=\frac{-p^2}{\hbar^2}\psi(x,t)"


"E\psi(x,t)=\frac{p^2}{2m}\psi(x,t)+V(x)\psi(x, t)\rightarrow(1)"

"P=\hbar k\\k=\frac{2\pi}{\lambda}"


"E\psi(x,t)=\frac{-\hbar^2}{2m}\frac{d^2}{dt^2}\psi(x,t)+V(x)\psi(x,t)"

"E=\hbar w"

"E\psi(x,t)=-\frac{\hbar w}{iw}\psi(x,t)"

"i\hbar\frac{d\psi}{dt}=-\frac{\hbar^2}{2m}\frac{d^2}{dx^2}\psi(x,t)+V(x)\psi(x,t)"